In Desmos, type the equation: f(x)= 1/2 square root of x-1


  1. What are the transformations given the equation?
1
Expert's answer
2021-03-18T15:50:12-0400

"g(x)=\sqrt x" red function

"f(x) =\frac{1}{2}\sqrt{x-1}" blue function




Transformations:

"\sqrt x\to \frac{1}{2}\sqrt x" multiplication by "\frac{1}{2}" shrinks the graph in half

"\frac{1}{2}\sqrt x \to \frac{1}{2}\sqrt {x -1}" "–1" under the root shifts the graph to the right